An insurance policy pays 100 per day up to 3 days of hospitalization and 50 per day for each day of hospitalization thereafter. The number of days of hospitalization, X, is a discrete random variable with probability function ��=� ={ policy.

1
Expert's answer
2021-11-30T06:13:42-0500

Letting "X" denote the number of days in the hospital, and "Y" the total payment, the given information can be summarized in the following table: 


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 k & 1 & 2 & 3 & 4 & 5 \\ \hline
 P(X=k) & 5/15 & 4/15 &3/15 & 2/15 & 1/15 \\
 \hdashline
 Y & 100 & 200 & 300 & 350 & 400 \\
 
\end{array}"


"P(Y)=100(\dfrac{5}{15})+200(\dfrac{4}{15})+300(\dfrac{3}{15})"

"+350(\dfrac{2}{15})+400(\dfrac{1}{15})=220"

The the expected payment for hospitalization under this policy is 220.